# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ 1.2 Contributions of the Paper
As a special instance in our framework, the FAIR neural network (FAIR-NN) estimator is proposed for which $\mathcal{G}$ and $\mathcal{F}$ are neural network classes to unveil $m^{\star}$ in the above nonparametric invariance pursuit problem. To the best of our knowledge, it is the first theoretically guaranteed estimator that can efficiently recover $m^{\star}$ under a single general and minimal identification condition associated with the heterogeneity of the environments. Specifically, the sample efficiency of FAIR-NN can be understood in several notable aspects: it requires the minimal identification condition, leading to fewer required environments; it exhibits the same $L_{2}$ error rate as if directly regressing $Y$ on known $X_{S^{\star}}$, regardless of the complexity of spurious associations; and it adapts to the unknown low-dimension structure of the invariant association $m^{\star}$ in a same manner as Kohler & Langer, (2021). In a nutshell, the FAIR-NN estimator circumvents the “curse of dimensionality” and “curse of endogeneity” simultaneously in a completely algorithmic manner, which does not rely on the prior knowledge of $m^{\star}$ structure or cause-effect relationships among variables. After attaining statistical estimation for the problem (1.1), we also exactly characterize the causal interpretation of $S^{\star}$ in the SCM with interventions setting under nondegenerate arbitrary interventions. This has further implications in robust transfer learning. A sufficient and necessary condition under which the direct causes of $Y$ can be recovered in general is also derived. These results are, to the best of our knowledge, the first in the literature and are of independent interest. This closes the crucial gap underscored in Section 1.1 and further affirms a positive answer to the previously posed Question Q to some extent.
@@
At the methodology level, our method bridges the invariance framework (Peters et al.,, 2016) and asymmetry principle (Janzing et al.,, 2016) for observational data into a unified framework. Besides using the invariance structure to identify $m^{\star}$, our framework also allows the users to incorporate their strong prior structural knowledge such as linearity or additivity of $m^{\star}$ into the FAIR estimation. This can be realized by restricting the function class $\mathcal{G}$ within this known structure and designating $\mathcal{F}$ as a more expansive class. We also propose the FAIR-AugLinear ($\mathcal{G}$ is linear and $\mathcal{F}$ is linear augmented with one transformed feature) and FAIR-ANN ($\mathcal{G}$ is additive neural network, $\mathcal{F}$ is neural network) estimators in Appendix B which leverage the prior knowledge that $m^{\star}$ is linear or additive, respectively. We demonstrate that harnessing such strong structural knowledge can relax the condition for identification. It is worth pointing out that identification is viable even when $|\mathcal{E}|=1$ for the FAIR-AugLinear estimator in Section B.6.
1.3 Related Works
@@
Causal Invariance and Invariant Risk Minimization. Starting from the pioneering work of Peters et al., (2016), there is considerable literature proposing methods to estimate $m^{\star}$, predominantly when $m^{\star}$ is linear. These methods broadly fall into two categories: hypothesis test-based methods and optimization-based methods. For the hypothesis test-based methods (Peters et al.,, 2016; Heinze-Deml et al.,, 2018; Pfister et al.,, 2019), the Type-I error is controlled for an estimator $\widehat{S}$ with $\mathbb{P}(\widehat{S}\subseteq S^{\star})\geq 1-\alpha$. Nonetheless, these procedures may result in missing important variables or conservative solutions like $\widehat{S}=\emptyset$ due to the inherent worst-case construction in the algorithm. Additionally, the introduction of hypothesis tests also hinders its seamless integration by machine learning algorithms, limiting their scalability. On the other hand, some optimization-based methods (Ghassami et al.,, 2017; Rothenhäusler et al.,, 2019, 2021) focus on linear $m^{\star}$ and tackle the problem under additional structures such as linear SCMs with additive interventions (Rothenhäusler et al.,, 2019). The additional structure imposed curtails its applicability to a broader nonparametric setting. There are also optimization-based methods (Pfister et al.,, 2021; Yin et al.,, 2021) designed for linear models, but they are heuristic and lack finite sample guarantees. In summary, there is still a crucial gap towards efficiently estimating $m^{\star}$ without additional assumptions on the underlying model. Although Fan et al., (2023) recently bridged this gap for linear $m^{\star}$ through an optimization-based method, it is still unclear under the general nonparametric setting. This paper is the first to attain sample-efficient estimation for the general model with non-asymptotic guarantees in terms of both $|\mathcal{E}|$ and $n$.
@@
Arjovsky et al., (2019) considers a general task, which aims to search for a data representation such that the optimal solution given that representation is optimal across diverse environments. They propose an optimization-based approach called invariant risk minimization (IRM), with many subsequent variants proposed later. However, their method requires at least $d$ environments even for the linear model, and the improvement over standard empirical risk minimization is not clear (Rosenfeld et al.,, 2021; Kamath et al.,, 2021). Our paper is the first to offer a comprehensive theoretical analysis of general invariance learning when the representation class is $\{(x_{1},\ldots,x_{d})\to(a_{1}x_{1},\ldots,a_{d}x_{d}):a_{1},\ldots,a_{d}\in\{0,1\}\}$ and to show that sample efficient estimation is in general viable even when $|\mathcal{E}|=2$.
@@
Causal Discovery without Active Intervention. There is considerable literature on causal discovery using observational data (Spirtes et al.,, 2000; Richardson,, 1996; Chickering,, 2002; Hyttinen et al.,, 2013, 2014), but they only work on the Markov equivalence class (Meek,, 1995; Geiger & Pearl,, 1990), and cannot identify the specific cause-effect relationship in general. Under the SCM framework, there is also considerable literature capable of determining the cause-effect direction under additional assumptions imposed under the circumstances that the algorithm can only passively observe data rather than performing intervention actively. These methods can be divided into two categories – one based on the invariance principle and one based on the asymmetry principle. The invariance-based approaches (Peters et al.,, 2016) consider collecting samples from multiple experiments where some unknown intervention may apply to the variables other than $Y$. It leverages the idea that the cause-effect mechanism will remain consistent while the reverse effect-cause association may vary to identify the cause-effect relationship. On the other hand, the asymmetry-based approaches (Shimizu et al.,, 2006; Hoyer et al.,, 2008; Zhang & Hyvärinen,, 2009; Janzing et al.,, 2012; Peters et al.,, 2014) only observe one sample of observational data and identify the cause-effect direction using the idea that the cause-effect mechanism admits a simple prior known structure, whereas its inverse does not, example includes the additive noise structure in Hoyer et al., (2008). These two principles for causal discovery seem to have been orthogonal before. To the best of our knowledge, our estimation framework is the first to offer a unified methodological perspective on these two principles. It demonstrates the ability to simultaneously leverage both principles for identification and estimation; see Section B.6.
@@ 1.4 Organization
The paper is structured as follows. We provide a generalized invariance pursuit problem in Section 2 and then present our unified estimation framework in Section 3. The main text afterward mainly focuses on the nonparametric invariance pursuit in (1.1), a special instance of our framework, to illustrate the main idea and philosophy: we show in Section 4.1 that applying our framework using neural networks can attain adaptive sample-efficient estimation therein, and show that $m^{\star}$ in (1.1) can be interpreted as a type of data-driven causality via characterizing the semantic meaning of $S^{\star}$ under the SCM with intervention setting. We provide a sketch of the abstract unified theoretical result, from which all non-asymptotic results are derived as corollaries, along with its other applications in Section 5 and defer the detailed statements to the appendix. Several extensions of our framework are discussed in Section A.3. We provide a computationally efficient implementation using variants of gradient descent and Gumbel approximation, followed by its application to the simulation and real data analysis in Section 6. All the proofs are collected in the supplemental material.
@@ 1.5 Notations
In the context of the multi-environment setup, consider the following notations. For each $e\in\mathcal{E}$, let $\Theta^{(e)}=L_{2}(\mu^{(e)}_{x})$ be the set of $L_{2}$ functions with respect to $\mu_{x}^{(e)}$, i.e., $\Theta^{(e)}=\{f:\int f^{2}(x)\mu_{x}^{(e)}(dx)<\infty\}$, and denote $\|f\|_{2,e}=\{\int f^{2}(x)\mu^{(e)}_{x}(dx)\}^{1/2}$. Moreover, suppose we have $n$ observations $(X_{1}^{(e)},Y_{1}^{(e)}),\ldots,(X_{n}^{(e)},Y_{n}^{(e)})\in\mathbb{R}^{d}\times\mathbb{R}$ drawn i.i.d. from $\mu^{(e)}$, we define $\mathbb{E}[f(X^{(e)},Y^{(e)})]=\int f(x,y)\mu^{(e)}(dx,dy)$ and $\widehat{\mathbb{E}}[f(X^{(e)},Y^{(e)})]=\frac{1}{n}\sum_{i=1}^{n}f(X^{(e)}_{i},Y^{(e)}_{i})$ for any $f\in\Theta^{(e)}$. We define $\bar{\mu}=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}$, and $\Theta=L_{2}(\bar{\mu}_{x})$ equipped with the norm $\|\cdot\|_{2}=\{\int f^{2}(x)\bar{\mu}_{x}(dx)\}^{1/2}$. It is easy to verify that $\Theta=\bigcap_{e\in\mathcal{E}}\Theta^{(e)}$. We assume $\mathbb{E}[|Y^{(e)}|^{2}]<\infty$.
@@
Neural Networks. We will use neural networks as a scalable nonparametric technique: we adopt the fully connected deep neural network with ReLU activation $\sigma(\cdot)=\max\{0,\cdot\}$, and call it deep ReLU network for short. Let $L,N$ be any positive integer, a deep ReLU network with depth $L$ width $N$ admits the form of

### v2

@@ 1.3 New Contributions
As a special instance in our framework, the FAIR neural network (FAIR-NN) estimator is proposed for which $\mathcal{G}$ and $\mathcal{F}$ are neural networks to unveil $m^{\star}$ in (1.1). It is the first theoretically guaranteed estimator that can efficiently recover $m^{\star}$ under a single general and minimal identification condition associated with the heterogeneity of the environments. Its sample efficiency can be understood in several notable aspects: it requires the minimal identification condition, leading to fewer required environments; it exhibits the same $L_{2}$ error rate as if directly regressing $Y$ on known $X_{S^{\star}}$, regardless of the complexity of spurious associations; and it adapts to the unknown low-dimension structure of the invariant association $m^{\star}$ in a same manner as Kohler & Langer, (2021). In summary, the FAIR-NN estimator circumvents the “curse of dimensionality” and “curse of endogeneity” simultaneously in a fully algorithmic manner, which does not rely on the prior knowledge of $m^{\star}$ structure or cause-effect relationships among variables.
While the complicated combinatorial constraint and minimax optimization are introduced in (1.2), we show that a variant of gradient descent – gradient descent ascent with Gumbel approximation to handle the combinatorial-nature “focused” constraint $f\in\mathcal{F}_{S_{g}}$ – continues to apply to our specifically designed algorithm and neural network estimators with no curse-of-dimension in implementation. Numerical results in Section 5 support this.
@@
Though our framework is designed for algorithmic learning, it is versatile in that the user can also incorporate their strong prior structural knowledge such as linearity or additivity of $m^{\star}$ into the FAIR estimation. This can be realized by restricting the function class $\mathcal{G}$ within this known structure and designating $\mathcal{F}$ as a more expansive class. We demonstrate that harnessing such strong structural knowledge can relax the condition for identification. It is worth pointing out that identification is viable even when $|\mathcal{E}|=1$ corresponding to observational data; see examples in Section B.6. At the methodology level, our method bridges the invariance principle (Peters et al.,, 2016) and asymmetry principle (Janzing et al.,, 2016) for observational data into a unified framework.
1.4 Related Works and Comparisons
@@
Starting from the pioneering work of Peters et al., (2016), there is considerable literature proposing methods to estimate $m^{\star}$ in (1.1), predominantly when $m^{\star}$ is linear. These methods broadly fall into two categories: hypothesis test-based methods and optimization-based methods. For the hypothesis test-based methods (Peters et al.,, 2016; Heinze-Deml et al.,, 2018; Pfister et al.,, 2019), the Type-I error is controlled for an estimator $\widehat{S}$ with $\mathbb{P}(\widehat{S}\subseteq S^{\star})\geq 1-\alpha$. Nonetheless, these procedures may result in missing important variables or conservative solutions like $\widehat{S}=\emptyset$ due to the inherent worst-case construction in the algorithm. Additionally, the introduction of hypothesis tests also hinders its seamless integration by machine learning algorithms, limiting their scalability. On the other hand, some optimization-based methods (Ghassami et al.,, 2017; Rothenhäusler et al.,, 2019, 2021) focus on linear $m^{\star}$ and tackle the problem under additional structures such as linear SCMs with additive interventions (Rothenhäusler et al.,, 2019). This limitation curtails its applicability to a broader nonparametric setting. Some optimization-based methods (Pfister et al.,, 2021; Yin et al.,, 2021) designed for linear models are heuristic and lack finite sample guarantees. In summary, there is still a crucial gap towards efficiently estimating $m^{\star}$ without additional assumptions on the underlying model. Although Fan et al., (2023) recently bridged this gap for linear $m^{\star}$ through an optimization-based method, it is still unclear under the general nonparametric setting. This paper is the first to attain sample-efficient estimation for the general model with non-asymptotic guarantees in terms of both $|\mathcal{E}|$ and $n$.
@@
Arjovsky et al., (2019) considers a general task, which aims to search for a data representation such that the optimal solution given that representation is optimal across diverse environments. They propose an optimization-based approach called invariant risk minimization (IRM), with many subsequent variants proposed later. However, their method comes with no statistical guarantees and requires at least $d$ environments even for the linear model, and the improvement over standard empirical risk minimization is not clear (Rosenfeld et al.,, 2021; Kamath et al.,, 2021). Our paper is the first to offer a comprehensive theoretical analysis of general invariance learning when the representation class is $\{(x_{1},\ldots,x_{d})\to(a_{1}x_{1},\ldots,a_{d}x_{d}):a_{1},\ldots,a_{d}\in\{0,1\}\}$ and to show that sample efficient estimation is in general viable even when $|\mathcal{E}|=2$. The main reason why this is attainable is due to the exact invariance pursued by our FAIR penalty and its “focused” nature, see the discussion in Section A.2.
@@
Under the SCM framework, there is considerable literature on causal discovery using observational data (Spirtes et al.,, 2000; Richardson,, 1996; Chickering,, 2002; Hyttinen et al.,, 2013, 2014), but they cannot go beyond Markov equivalent class (Geiger & Pearl,, 1990) and thus fail to establish the exact cause-effect direction in general. Such a problem can be resolved by imposing additional assumptions under the circumstances that the algorithm can only passively observe data rather than performing intervention actively. These methods can be divided into two categories – one based on the invariance principle and the other based on the asymmetry principle. The invariance-based approaches (Peters et al.,, 2016) use samples from multiple experiments where some unknown intervention may apply to the variables other than $Y$. It leverages the idea that the cause-effect mechanism will remain constant while the reverse effect-cause association may vary. On the other hand, the asymmetry-based approaches (Shimizu et al.,, 2006; Hoyer et al.,, 2008; Zhang & Hyvärinen,, 2009; Janzing et al.,, 2012; Peters et al.,, 2014) only observe one sample of observational data and use the idea that the cause-effect mechanism admits a simple prior known structure, whereas its inverse does not, example includes the additive noise structure (Hoyer et al.,, 2008). These two principles for causal discovery seem to have been orthogonal before. Our estimation framework is the first to offer a unified methodological perspective on these two principles with theoretical guarantees. It demonstrates the ability to simultaneously leverage both principles for identification and estimation.
Adversarial estimation is introduced in Goodfellow et al., (2014) for generative modeling. Its application in the statistics spans distribution estimation (Liang,, 2021), instrumental variable regression (Dikkala et al.,, 2020), estimating the (implicit) influence function (Chernozhukov et al.,, 2020; Hirshberg & Wager,, 2021), and so on. The idea of minimizing the worst-case reward among diverse environments can also be considered as “an algorithmic remedy” for out-of-distribution generalization. There are different considerations of the “reward” such as risk (Sagawa et al.,, 2020), excess risk (Agarwal & Zhang,, 2022), and the negative of the explained variance (Meinshausen & Bühlmann,, 2015). However, these methods are conceptually similar to running least squares in regression and thus cannot go beyond just learning associations. We adopted the adversarial estimation in our estimation from two novel aspects. Firstly, it allows us to use a simple objective function that homogenizes different tasks and prediction models for estimation. Moreover, such a minimax optimization objective and the Gumbel approximation in the implementation jointly relax the combinatorial nature in (1.3) and make a variant of gradient descents continue to work numerically.
@@ 1.5 Organization
This paper is structured as follows. We first provide the proposed method with non-asymptotic theoretical analysis, and causal interpretations for our canonical nonparametric causality (invariance) pursuit problem in Sections 23, respectively. Such a special instance of our framework also helps to illustrate the main idea and philosophy of our general invariance pursuit problem and FAIR estimation framework, which will be formally presented in Section 4. In the main text, we provide a sketch of the abstract unified result, from which all non-asymptotic results are derived as corollaries, along with its other applications in Section 4.3 and defer the detailed statements to the Appendix. We provide a computationally efficient implementation using variants of gradient descent and Gumbel approximation, followed by its application to the simulation and real data analysis in Section 5. All the proofs are collected in the supplemental material.
@@ 1.6 Notations
In the context of the multi-environment setup, consider the following notations. For each $e\in\mathcal{E}$, let $\Theta^{(e)}=L_{2}(\mu^{(e)}_{x}):=\{f:\int f^{2}(x)\mu_{x}^{(e)}(dx)<\infty\}$, and denote $\|f\|_{2,e}=\{\int f^{2}(x)\mu^{(e)}_{x}(dx)\}^{1/2}$. Given $n$ observations $\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ drawn i.i.d. from $\mu^{(e)}$, we define $\mathbb{E}[f(X^{(e)},Y^{(e)})]=\int f(x,y)\mu^{(e)}(dx,dy)$ and $\widehat{\mathbb{E}}[f(X^{(e)},Y^{(e)})]=\frac{1}{n}\sum_{i=1}^{n}f(X^{(e)}_{i},Y^{(e)}_{i})$ for any $f\in\Theta^{(e)}$. We assume $\mathbb{E}[|Y^{(e)}|^{2}]<\infty$. Let $\bar{\mu}=\frac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mu^{(e)}$, and $\Theta=L_{2}(\bar{\mu}_{x})$ equipped with the norm $\|\cdot\|_{2}=\{\int f^{2}(x)\bar{\mu}_{x}(dx)\}^{1/2}$. It is easy to verify that $\Theta=\bigcap_{e\in\mathcal{E}}\Theta^{(e)}$.
@@
Neural Networks. We use neural networks as a scalable nonparametric technique: we adopt the fully connected deep neural network with ReLU activation $\sigma(\cdot)=\max\{0,\cdot\}$, and call it deep ReLU network for short. Let $L,N$ be any positive integer, a deep ReLU network with depth $L$ width $N$ admits the form of